If a=3,b=4,c=5 then sin2B is....plzz I want an answer urgently!!
 

Grade:9

2 Answers

Param Prashar
8 years ago
a=3
b=4
c=5
32+42=52
therefore it is a pythagoran pair
\angleC=90o 
sinB=4/5
cosB=3/5
sin2B=2sinBcosB=2*(4/5)*3/5
Ankit Singh
7 years ago
Hey Drovan this question is done by the properties and solution of triangles tooHere in this problemGiven : a=3, b=4 and c=5To find : sin2BSolution : we know that sin2B = 2sinBcosBTo solve this problem we have to take angle B of triangle as right angle therefore , 2sinBcosB= 2(b/c)(c²+a²-b²/2ca) = 2[(4/5) (5²+3²-4²/2×5×3)] = 2[(4/5) (25+9-16/30)] = 2[(4/5) (18/30)] = 2×4/5×18/30 = 24/25 (Ans.)Since :- in right angle triangle sin B = b/c and cosB= c²+a²-b²/2caI hope you can like this if it is in respect with properties and solutions of triangles